In a double‑slit interference experiment, the wavelength is ?=532 nm , the slit separation is ?=0.190 mm , and the screen is ?=50.0 cm away from the slits. What is the linear distance Δ? between the seventh order maximum and the fifth order maximum on the screen?
Answer #1

condition for constructive interference is

d sin theta = m lamda

for small angles sin theta = tan theta = y/L

d ( y/D) = m lamda

for m = 7

y7= m lamda D/ d

= 7(532* 10^-9) (50* 10^-2)/0.190 * 10^-3 m

=9.8 * 10^-3 m

for m = 5

y5= m lamda D/ d

= 5(532* 10^-9) (50* 10^-2)/0.190 * 10^-3 m

=7 * 10^-3 m

del x = y7 - y5 = 9.8 * 10^-3 m -7* 10^-3 m = 2.8 * 10^-3 m = 2.8 mm